# ThinK: Thinner Key Cache by Query-Driven Pruning

### v1

@@ 3.2 Query-Driven Pruning
For each head, the attention scores are computed using the queries and keys, and then applied to the values. The formula for the attention for head $i$ is:
$\text{Attention}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})=\text{softmax}(\frac{\textbf{Q}_{i}\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}$ (5)
where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. When pruning one channels of $\textbf{K}_{i}$, the corresponding channels in $\textbf{Q}_{i}$ will also be removed. We aim to find the optimal subset of channels to prune, denoted by the selection matrix $\textbf{S}\in\{0,1\}^{D\times D}$, where S is a diagonal matrix with binary entries ($1$ for keeping a channel, 00 for pruning it). To better maintain the performance after pruning the channels, we minimize the Frobenius norm of the difference between the original and pruned attention weights:
$\min_{\textbf{S}}\|\textbf{Q}_{i}\textbf{K}^{T}_{i}-\textbf{Q}_{i}\textbf{S}(\textbf{K}_{i}\textbf{S})^{T}\|_{F}$ (6)
Given a pruning ratio $\lambda$, it can further expanded as:
$\min_{\mathbf{S}}$ $\quad\left\|\mathbf{Q}_{i}\mathbf{K}_{i}^{T}-\mathbf{Q}_{i}\mathbf{S}\mathbf{K}_{i}^{T}\right\|_{F}$
@@
$\quad\mathbf{S}=\text{diag}(s_{1},s_{2},\ldots,s_{D}),\text{ where }s_{j}\in\{0,1\}$ (7)
For simplicity, we use greedy algorithm to optimize $\mathbf{S}$. To achieve the pruning goal, we define a criterion for evaluating the importance of each channel and greedily select the channels with largest scores:
$\text{Score}_{i}[j]=\big{\|}\textbf{Q}_{i}[:,j]\textbf{K}_{i}[:,j]^{T}\big{\|}_{F}$ (8)
$I_{i}=\textbf{Top}_{T}(\text{Score}_{i},T)$ (9)
Here’s a detailed explanation of why it optimizes the selection matrix. The $\text{score}_{i}[j]$ measures the magnitude of the interaction between the query and key vectors for channel $j$ in each head $i$. By selecting channels with the highest interaction magnitudes, we aim to retain the most significant contributions to the attention mechanism. This criterion ensures that the selected channels preserve the primary information flow in the attention computation, thereby minimizing the loss of important information.
Observation Window. Based on the observations in SnapKV (Li et al.,, 2024) that the last window of input sequence recognizes highly similar attention allocation pattern with the actual generation. To reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
@@
3.3 Implementations
Following SnapKV, we focus on the long context input scenario. We opt not to prune the most recent tokens and newly generated keys. Consequently, our key-value (KV) cache will store two distinct categories of keys: one subset consists of pruned keys with a reduced channel size, while the other retains keys at their original size. Additionally, we maintain a binary mask to indicate which channels have been pruned. Note that the memory overhead associated with this mask is negligible. Figure 4 illustrates one implementation of our method during the decoding stage. The pruned keys are first zero-filled using the mask to restore them to their original size, and then concatenated with the unpruned keys. This implementation can be effectively integrated with optimization techniques such as FlashAttention (Dao,, 2023). The second implementation involves initially pruning the Query using the mask. The pruned Query is then multiplied by the pruned Key, while the unpruned Query is applied to the unpruned Key. Subsequently, the two outputs are concatenated. In theory, this implementation has lower computational costs.
4 Experiment Results
In this section, we conduct extensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction. First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1. Next, we report the performance of our KV pruning method with two different KV compression methods in long-context scenarios in Section 4.2. Finally, we present the results of ablation experiments in Section 4.4.
4.1 Settings
Benchmark Datasets.
We perform evaluations of our method with state-of-the-art KV cache compression methods on two widely used benchmarks: LongBench and Needle-in-a-Haystack. LongBench (Bai et al.,, 2023) is designed to comprehensively evaluate LLM’s long context understanding capabilities. It includes 17 datasets covering six different tasks: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. The average input length of LongBench is 6,711 words, which necessitates reducing the KV cache to lower memory usage for inference. Needle-in-a-Haystack (Kamradt,, 2023) is a recent popular test challenge that requires models to accurately identify a small piece of information (“needle”) in a long document (“haystack”), where the needle is placed at a random position. This challenge can test if KV cache compression methods still retain the small piece of critical information.
Baseline Approaches.
The baseline methods include Heavy Hitter Oracle (H2O) and SnapKV, both of which are the state-of-the-art KV cache compression methods but use different strategies for selecting KV cache. H2O (Zhang et al., 2024c, ) is designed to reduce memory usage by dynamically balancing recent tokens and Heavy Hitter (H2) tokens, where H2 tokens are a small set of tokens that contribute most of the value when computing attention scores. SnapKV (Li et al.,, 2024) automatically compresses KV caches by selecting clustered important KV positions for each attention head. This is inspired by the fact that each attention head has a specific pattern, focusing on certain attention features during generation.
Table 2: Performance comparison of Key cache pruning on LLaMA3-8B by ThinK on LongBench.
@@ 4.1 Settings
KV-size Full
@@
+ThinK(40) 22.11 14.67 32.49 36.25 28.63 18.80 18.93 21.49 20.14 44.50 88.11 38.32 5.75 69.17 58.21 55.89 35.84
+ThinK(50) 21.79 14.73 32.03 37.52 27.86 18.28 18.50 21.52 19.71 43.50 86.00 38.35 5.59 69.50 57.96 56.96 35.61
KV-size 512
@@
+ThinK(40) 23.76 17.80 33.80 40.39 30.70 19.09 21.82 22.51 23.78 41.00 90.16 40.67 5.15 69.25 60.77 57.58 37.39
+ThinK(50) 24.17 16.96 35.76 39.47 30.29 18.67 21.39 22.59 23.06 41.00 89.81 40.35 5.23 69.33 60.20 58.34 37.29
+ThinK(60) 23.40 14.83 32.62 38.47 30.97 19.81 20.80 22.04 21.60 40.00 88.79 38.90 5.36 69.50 58.28 57.65 36.44
@@
+ThinK(40) 24.58 25.44 37.03 41.87 33.45 20.58 21.77 22.42 24.16 70.00 90.39 40.29 6.06 69.50 62.05 59.23 40.55
+ThinK(50) 24.85 25.10 37.06 41.58 32.34 20.60 21.61 22.44 23.66 69.50 90.39 39.70 5.84 69.79 61.57 59.42 40.34
+ThinK(60) 25.98 22.77 38.37 40.44 33.19 19.90 20.84 22.21 22.55 59.00 90.32 38.12 6.39 69.50 59.14 58.40 39.20
KV-size 1024
@@
+ThinK(40) 25.52 21.93 37.17 41.56 31.22 20.17 22.89 22.95 25.21 47.00 90.74 41.34 5.57 69.50 62.58 58.67 39.00
+ThinK(50) 25.41 22.19 37.64 40.92 31.27 18.66 22.17 22.22 24.84 46.50 90.34 40.59 5.20 69.50 61.71 57.99 38.57
+ThinK(60) 24.06 17.80 37.85 38.63 29.98 19.40 21.41 22.32 23.42 44.50 90.16 39.43 5.84 69.50 58.31 58.73 37.58
@@
+ThinK(40) 24.88 27.72 38.60 43.16 32.44 20.67 24.21 22.79 25.56 71.50 90.45 40.94 5.93 69.50 62.77 59.45 41.29
+ThinK(50) 24.82 27.26 39.66 42.82 32.09 19.56 23.52 22.48 25.34 71.50 90.43 40.74 5.20 69.50 62.46 59.75 41.07
+ThinK(60) 24.46 27.35 38.22 41.96 31.64 20.18 21.89 22.83 23.68 70.00 90.19 38.69 6.10 69.50 58.87 59.26 40.30
KV-size 2048
@@
+ThinK(40) 25.56 26.31 39.20 42.96 31.81 20.53 24.23 23.35 25.90 53.50 90.56 41.03 5.52 69.25 62.10 59.00 40.05
+ThinK(50) 25.01 25.37 38.82 42.32 31.27 20.50 23.78 23.21 26.03 53.00 90.37 40.86 5.13 69.50 61.91 58.95 39.75
+ThinK(60) 24.37 22.14 37.77 40.13 29.50 20.26 22.09 22.76 24.78 49.50 90.16 39.69 5.56 69.50 59.24 58.78 38.51
@@
+ThinK(40) 25.41 29.79 39.21 43.35 33.96 21.49 25.78 23.11 26.23 73.00 90.56 41.79 5.81 69.50 62.45 59.19 41.91
+ThinK(50) 25.00 30.25 39.27 43.23 32.93 21.24 25.16 23.01 26.5 73.00 90.37 41.26 5.45 69.50 62.3 59.84 41.77
+ThinK(60) 24.89 28.88 40.44 41.30 29.99 21.34 23.48 22.9 24.99 72.50 90.36 38.5 5.71 69.50 59.77 59.50 40.88
Implementation Details.
In this paper, we use LLaMA3-8B-Instruct (Meta,, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al.,, 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al.,, 2020). Our ThinK aims to prune channels of the KV, which is agnostic to KV cache compression methods. If there is no other statement, we prune the key cache by default. All the experiments are conducted using one Nvidia A100. To fairly compare KV cache compression methods and KV cache compression integrated with ThinK, we use the same hyperparameters for both. For example, when comparing SnapKV and SnapKV integrated with ThinK, we set the maximum pooling kernel size to 7 and the observation window size to 32, using the same KV-size for both.
Table 3: Performance comparison of Key cache pruning on Mistral-7B by ThinK on LongBench.
@@
KV-size Full
@@
+ThinK(40) 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
+ThinK(50) 21.67 21.80 39.48 28.74 20.65 13.14 20.57 22.83 21.78 39.00 82.54 40.12 3.61 78.39 50.27 48.4 34.56
+ThinK(60) 21.04 21.30 39.56 28.68 21.29 13.97 20.13 22.52 21.81 39.50 82.05 39.14 4.16 74.23 49.83 47.67 34.18
@@
+ThinK(40) 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
+ThinK(50) 20.67 20.60 43.37 37.27 21.58 15.66 19.06 21.79 21.02 47.00 83.38 39.77 3.65 67.06 50.80 48.35 35.06
+ThinK(60) 21.25 20.82 44.20 36.21 21.68 16.47 19.05 21.99 20.73 45.00 83.81 38.79 4.19 66.90 49.99 47.61 34.92
KV-size 512
@@
+ThinK(40) 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
+ThinK(50) 22.76 25.74 44.61 31.74 23.25 13.91 23.31 23.13 24.34 41.00 85.39 41.85 2.82 84.36 54.69 50.88 37.11
+ThinK(60) 22.91 25.57 44.04 29.48 22.88 13.67 23.31 22.64 24.10 41.00 85.31 41.15 2.98 82.34 53.70 50.25 36.58
@@
+ThinK(40) 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.9 42.51 2.92 85.32 55.89 53.35 40.40
+ThinK(50) 24.56 29.22 48.59 37.70 24.27 17.39 23.68 23.65 24.58 67.50 86.05 42.01 3.07 86.30 56.49 53.29 40.52
+ThinK(60) 24.07 28.27 49.10 38.65 24.31 17.52 23.16 23.51 24.23 67.00 86.33 40.78 3.69 83.74 54.94 52.23 40.10
KV-size 1024
@@
+ThinK(40) 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
+ThinK(50) 23.89 28.40 46.60 35.57 24.26 14.78 24.98 23.31 25.68 44.50 86.16 42.72 3.38 83.20 55.88 52.63 38.50
+ThinK(60) 23.87 27.76 46.25 35.28 24.38 14.74 24.35 23.35 25.50 44.50 85.38 41.37 3.34 81.42 55.21 51.89 38.04
@@
+ThinK(40) 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.3 56.71 54.19 41.62
+ThinK(50) 25.63 30.08 49.41 40.59 25.13 19.58 25.47 24.23 25.92 69.5 86.67 42.31 2.74 84.78 57.43 53.59 41.44
+ThinK(60) 24.69 29.3 48.90 40.44 25.33 19.58 25.23 23.6 25.25 69.00 86.85 40.86 3.19 83.70 56.3 53.30 40.97
KV-size 2048
@@
+ThinK(40) 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
+ThinK(50) 25.68 31.24 48.69 39.65 25.84 16.72 26.69 23.57 26.78 52.00 86.74 42.85 4.01 83.46 57.12 53.67 40.29
+ThinK(60) 25.83 31.00 48.23 38.58 25.71 16.54 26.51 23.81 26.28 50.50 86.57 42.05 3.36 82.49 56.04 52.67 39.76
@@
+ThinK(40) 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
+ThinK(50) 26.44 32.94 49.02 40.86 26.84 19.49 28.46 24.51 26.72 70.00 86.50 41.75 2.78 84.70 56.47 54.15 41.98
+ThinK(60) 26.00 32.53 48.73 40.95 26.77 18.92 27.40 23.97 26.37 70.00 86.45 41.12 3.31 82.24 56.01 53.53 41.52

### v2

@@ 3.2 Query-Driven Pruning
For each head, the attention scores are computed using the queries and keys, and then applied to the values. The formula for the attention for head $i$ is: $\text{Attention}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})=\text{softmax}(\frac{\textbf{Q}_{i}\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i},$ where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. When one channel of $\textbf{K}_{i}$ is pruned, the corresponding channel in $\textbf{Q}_{i}$ will also be removed. We aim to find the optimal subset of channels to prune, denoted by the selection matrix $\textbf{S}\in\{0,1\}^{D\times D}$, where S is a diagonal matrix with binary entries ($1$ for keeping a channel, 00 for pruning it). To better maintain the performance after pruning the channels, we minimize the Frobenius norm of the difference between the original and pruned attention weights: $\min_{\textbf{S}}\|\textbf{Q}_{i}\textbf{K}^{T}_{i}-\textbf{Q}_{i}\textbf{S}(\textbf{K}_{i}\textbf{S})^{T}\|_{F}.$ Given a pruning ratio $\lambda$, it can further expanded as:
$\min_{\mathbf{S}}$ $\quad\left\|\mathbf{Q}_{i}\mathbf{K}_{i}^{T}-\mathbf{Q}_{i}\mathbf{S}\mathbf{K}_{i}^{T}\right\|_{F}$ (2)
@@
$\quad\mathbf{S}=\text{diag}(s_{1},s_{2},\ldots,s_{D}),\text{ where }s_{j}\in\{0,1\}$
For simplicity, we use greedy algorithm to optimize $\mathbf{S}$. To achieve the pruning goal, we define a criterion for evaluating the importance of each channel and greedily select the channels with largest scores: $\text{Score}_{i}[j]=\big{\|}\textbf{Q}_{i}[:,j]\textbf{K}_{i}[:,j]^{T}\big{\|}_{F},\quad I_{i}=\textbf{Top}_{T}(\text{Score}_{i},T)$. Here’s a detailed explanation of why it optimizes the selection matrix. The $\text{score}_{i}[j]$ measures the magnitude of the interaction between the query and key vectors for channel $j$ in each head $i$. By selecting channels with the highest interaction magnitudes, we aim to retain the most significant contributions to the attention mechanism. This criterion ensures that the selected channels preserve the primary information flow in the attention computation, thereby minimizing the loss of important information.
Observation Window. Based on the observations in SnapKV (Li et al., 2024) that the last window of input sequence recognizes highly similar attention pattern with generation. To reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
3.3 Implementation of ThinK
@@
Following SnapKV, we focus on the long context input scenario. We opt not to prune the most recent tokens and newly generated keys. Consequently, our key-value (KV) cache will store two distinct categories of keys: one subset consists of pruned keys with a reduced channel size, while the other retains keys at their original size. Additionally, we maintain a binary mask to indicate which channels have been pruned. it is important to note that the memory overhead associated with this mask is negligible. Figure 2 illustrates one implementation of ThinK during the decoding stage. This implementation begins by pruning the query using the mask. The pruned query is then multiplied by the pruned key, while the unpruned query is applied to the unpruned Key. Subsequently, the two outputs are concatenated.
Table 2: Performance comparison of key cache pruning on LLaMA-3-(8B/70B)-Instruct on LongBench. ThinK ($\lambda$) indicates we prune the key cache channels with a pruning ratio of $\lambda$.
@@
LLaMA-3-8B-Instruct, KV-size Full
@@
+ThinK (0.4) 22.11 14.67 32.49 36.25 28.63 18.80 18.93 21.49 20.14 44.50 88.11 38.32 5.75 69.17 58.21 55.89 35.84
+ThinK (0.5) 21.79 14.73 32.03 37.52 27.86 18.28 18.50 21.52 19.71 43.50 86.00 38.35 5.59 69.50 57.96 56.96 35.61
LLaMA-3-8B-Instruct, KV-size 512
@@
+ThinK (0.4) 23.76 17.80 33.80 40.39 30.70 19.09 21.82 22.51 23.78 41.00 90.16 40.67 5.15 69.25 60.77 57.58 37.39
+ThinK (0.5) 24.17 16.96 35.76 39.47 30.29 18.67 21.39 22.59 23.06 41.00 89.81 40.35 5.23 69.33 60.20 58.34 37.29
+ThinK (0.6) 23.40 14.83 32.62 38.47 30.97 19.81 20.80 22.04 21.60 40.00 88.79 38.90 5.36 69.50 58.28 57.65 36.44
@@
+ThinK (0.4) 24.58 25.44 37.03 41.87 33.45 20.58 21.77 22.42 24.16 70.00 90.39 40.29 6.06 69.50 62.05 59.23 40.55
+ThinK (0.5) 24.85 25.10 37.06 41.58 32.34 20.60 21.61 22.44 23.66 69.50 90.39 39.70 5.84 69.79 61.57 59.42 40.34
+ThinK (0.6) 25.98 22.77 38.37 40.44 33.19 19.90 20.84 22.21 22.55 59.00 90.32 38.12 6.39 69.50 59.14 58.40 39.20
LLaMA-3-8B-Instruct, KV-size 1024
@@
+ThinK (0.4) 25.52 21.93 37.17 41.56 31.22 20.17 22.89 22.95 25.21 47.00 90.74 41.34 5.57 69.50 62.58 58.67 39.00
+ThinK (0.5) 25.41 22.19 37.64 40.92 31.27 18.66 22.17 22.22 24.84 46.50 90.34 40.59 5.20 69.50 61.71 57.99 38.57
+ThinK (0.6) 24.06 17.80 37.85 38.63 29.98 19.40 21.41 22.32 23.42 44.50 90.16 39.43 5.84 69.50 58.31 58.73 37.58
@@
+ThinK (0.4) 24.88 27.72 38.60 43.16 32.44 20.67 24.21 22.79 25.56 71.50 90.45 40.94 5.93 69.50 62.77 59.45 41.29
+ThinK (0.5) 24.82 27.26 39.66 42.82 32.09 19.56 23.52 22.48 25.34 71.50 90.43 40.74 5.20 69.50 62.46 59.75 41.07
+ThinK (0.6) 24.46 27.35 38.22 41.96 31.64 20.18 21.89 22.83 23.68 70.00 90.19 38.69 6.10 69.50 58.87 59.26 40.30
LLaMA-3-8B-Instruct, KV-size 2048
@@
+ThinK (0.4) 25.56 26.31 39.20 42.96 31.81 20.53 24.23 23.35 25.90 53.50 90.56 41.03 5.52 69.25 62.10 59.00 40.05
+ThinK (0.5) 25.01 25.37 38.82 42.32 31.27 20.50 23.78 23.21 26.03 53.00 90.37 40.86 5.13 69.50 61.91 58.95 39.75
+ThinK (0.6) 24.37 22.14 37.77 40.13 29.50 20.26 22.09 22.76 24.78 49.50 90.16 39.69 5.56 69.50 59.24 58.78 38.51
@@
+ThinK (0.4) 25.41 29.79 39.21 43.35 33.96 21.49 25.78 23.11 26.23 73.00 90.56 41.79 5.81 69.50 62.45 59.19 41.91
+ThinK (0.5) 25.00 30.25 39.27 43.23 32.93 21.24 25.16 23.01 26.5 73.00 90.37 41.26 5.45 69.50 62.3 59.84 41.77
+ThinK (0.6) 24.89 28.88 40.44 41.30 29.99 21.34 23.48 22.9 24.99 72.50 90.36 38.5 5.71 69.50 59.77 59.50 40.88
LLaMA-3-70B-Instruct, KV-size 128
SnapKV 25.91 39.41 43.83 49.60 51.23 27.76 22.14 21.91 23.16 69.00 91.55 43.54 12.50 72.00 48.41 63.49 44.09
+ThinK (0.4) 25.64 39.20 43.60 50.22 50.50 29.32 21.70 21.96 23.35 68.00 91.27 43.24 12.50 73.00 48.01 62.43 44.00
+ThinK (0.5) 26.31 38.76 44.86 48.54 49.62 28.97 21.46 22.01 22.91 67.00 91.52 43.15 12.50 72.50 47.21 60.82 43.63
LLaMA-3-70B-Instruct, KV-size 512
SnapKV 27.95 45.19 48.50 50.97 54.53 29.78 25.34 22.36 26.03 73.50 92.63 45.07 12.50 72.50 45.21 68.22 46.27
+ThinK (0.4) 27.47 45.31 48.57 51.22 54.32 30.05 25.42 22.72 26.20 73.50 92.13 45.53 12.50 73.00 48.32 66.99 46.45
+ThinK (0.5) 26.97 44.55 48.16 50.84 53.80 30.57 25.29 22.65 25.53 73.00 92.13 43.66 12.50 73.00 50.52 64.82 46.12
LLaMA-3-70B-Instruct, KV-size 1024
SnapKV 26.80 46.21 49.93 51.70 54.71 29.86 27.61 22.43 27.15 73.50 92.38 46.18 12.50 72.50 42.84 69.89 46.64
+ThinK (0.4) 27.04 46.01 50.13 51.96 54.36 29.87 27.74 22.78 27.07 73.50 91.88 46.35 12.50 73.00 45.05 67.87 46.69
+ThinK (0.5) 27.62 46.22 48.97 51.79 53.39 30.47 27.45 23.05 26.57 73.50 91.88 43.99 12.50 72.50 47.41 66.84 46.51
LLaMA-3-70B-Instruct, KV-size 2048
SnapKV 27.44 46.51 49.60 51.80 54.77 31.05 29.67 22.44 27.43 73.50 92.38 45.98 12.50 72.50 41.86 68.72 46.76
+ThinK (0.4) 27.13 46.26 50.04 51.72 55.03 31.19 29.75 22.47 27.28 73.50 91.88 46.37 12.50 72.50 42.66 67.77 46.75
+ThinK (0.5) 27.84 46.86 49.18 51.97 53.58 31.44 29.41 22.89 27.33 73.50 91.88 43.60 12.50 72.50 44.78 66.65 46.62
Table 3: Performance comparison of key cache pruning on Mistral-7B-Instruct-v0.2 on LongBench. ThinK ($\lambda$) indicates we prune the key cache channels with a pruning ratio of $\lambda$.
@@
Mistral-7B-Instruct-v0.2, KV-size Full
@@
+ThinK (0.4) 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
+ThinK (0.5) 21.67 21.80 39.48 28.74 20.65 13.14 20.57 22.83 21.78 39.00 82.54 40.12 3.61 78.39 50.27 48.4 34.56
+ThinK (0.6) 21.04 21.30 39.56 28.68 21.29 13.97 20.13 22.52 21.81 39.50 82.05 39.14 4.16 74.23 49.83 47.67 34.18
@@
+ThinK (0.4) 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
+ThinK (0.5) 20.67 20.60 43.37 37.27 21.58 15.66 19.06 21.79 21.02 47.00 83.38 39.77 3.65 67.06 50.80 48.35 35.06
+ThinK (0.6) 21.25 20.82 44.20 36.21 21.68 16.47 19.05 21.99 20.73 45.00 83.81 38.79 4.19 66.90 49.99 47.61 34.92
Mistral-7B-Instruct-v0.2, KV-size 512
@@
+ThinK (0.4) 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
+ThinK (0.5) 22.76 25.74 44.61 31.74 23.25 13.91 23.31 23.13 24.34 41.00 85.39 41.85 2.82 84.36 54.69 50.88 37.11
+ThinK (0.6) 22.91 25.57 44.04 29.48 22.88 13.67 23.31 22.64 24.10 41.00 85.31 41.15 2.98 82.34 53.70 50.25 36.58
@@
+ThinK (0.4) 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.9 42.51 2.92 85.32 55.89 53.35 40.40
+ThinK (0.5) 24.56 29.22 48.59 37.70 24.27 17.39 23.68 23.65 24.58 67.50 86.05 42.01 3.07 86.30 56.49 53.29 40.52
+ThinK (0.6) 24.07 28.27 49.10 38.65 24.31 17.52 23.16 23.51 24.23 67.00 86.33 40.78 3.69 83.74 54.94 52.23 40.10
Mistral-7B-Instruct-v0.2, KV-size 1024
@@
+ThinK (0.4) 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
+ThinK (0.5) 23.89 28.40 46.60 35.57 24.26 14.78 24.98 23.31 25.68 44.50 86.16 42.72 3.38 83.20 55.88 52.63 38.50
+ThinK (0.6) 23.87 27.76 46.25 35.28 24.38 14.74 24.35 23.35 25.50 44.50 85.38 41.37 3.34 81.42 55.21 51.89 38.04
@@
+ThinK (0.4) 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.3 56.71 54.19 41.62
+ThinK (0.5) 25.63 30.08 49.41 40.59 25.13 19.58 25.47 24.23 25.92 69.5 86.67 42.31 2.74 84.78 57.43 53.59 41.44
+ThinK (0.6) 24.69 29.3 48.90 40.44 25.33 19.58 25.23 23.6 25.25 69.00 86.85 40.86 3.19 83.70 56.3 53.30 40.97
Mistral-7B-Instruct-v0.2, KV-size 2048
@@
+ThinK (0.4) 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
+ThinK (0.5) 25.68 31.24 48.69 39.65 25.84 16.72 26.69 23.57 26.78 52.00 86.74 42.85 4.01 83.46 57.12 53.67 40.29
+ThinK (0.6) 25.83 31.00 48.23 38.58 25.71 16.54 26.51 23.81 26.28 50.50 86.57 42.05 3.36 82.49 56.04 52.67 39.76
@@
+ThinK (0.4) 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
+ThinK (0.5) 26.44 32.94 49.02 40.86 26.84 19.49 28.46 24.51 26.72 70.00 86.50 41.75 2.78 84.70 56.47 54.15 41.98
+ThinK (0.6) 26.00 32.53 48.73 40.95 26.77 18.92 27.40 23.97 26.37 70.00 86.45 41.12 3.31 82.24 56.01 53.53 41.52
4 Experiment Results
In this section, we conduct comprehensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction. First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1. Next, we report the performance ThinK with two KV eviction methods and one KV quantization methods in long-context scenarios in Section 4.2 and  4.3. Finally, we present the results of ablation experiments in Section 4.4.
4.1 Settings
Benchmark Datasets. We evaluate our proposed method against state-of-the-art KV cache compression methods on two widely recognized benchmarks: LongBench and Needle-in-a-Haystack. LongBench (Bai et al., 2023) is designed to comprehensively assess the long context understanding capabilities of LLMs. It includes 17 datasets covering six different tasks: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. The average input length of LongBench is 6,711 words, which necessitates reducing the KV cache to lower memory usage for inference. Needle-in-a-Haystack (Kamradt, 2023) is a recently developed benchmark that tests a model’s ability to accurately locate a small but crucial piece of information (the ”needle”) embedded within a lengthy document (the ”haystack”). The random positioning of the needle in this challenge serves as a critical test to determine whether KV cache compression methods can retain essential information without loss of accuracy.
Baseline Approaches. The baseline methods in our evaluations include Heavy Hitter Oracle (H2O), SnapKV and KIVI, all of which are the state-of-the-art KV cache compression methods but use different strategies. H2O (Zhang et al., 2024c) is designed to reduce memory usage by dynamically managing the balance between recent tokens and Heavy Hitter (H2) tokens. H2 tokens represent a small set of tokens that contribute most of the value when computing attention scores. SnapKV (Li et al., 2024) introduces an automated compression mechanism that selects clustered, important KV positions for each attention head, optimizing the KV cache without sacrificing performance. KIVI (Liu et al., 2024b) reduces memory overhead by quantizing the KV cache into lower-precision formats, significantly lowering the memory cost while preserving model accuracy.
Implementation Details. In this paper, we use LLaMA-2-7B-chat, LLaMA-3-8B-Instruct, LLaMA-3-70B-Instruct (Meta, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al., 2020). Our ThinK aims to prune channels of the key cache, which is agnostic to KV cache compression methods. If there is no other statement, we prune the key cache by default. All the experiments are conducted using NVIDIA A100 GPUs. To ensure a fair comparison between KV cache compression methods and their integration with ThinK, we applied consistent hyperparameters across both settings. For instance, when comparing SnapKV and SnapKV integrated with ThinK, we used a maximum pooling kernel size of 7 and an observation window size of 32, maintaining the same KV-size for both configurations.
Table 4: Performance evaluation of combining ThinK with KIVI Liu et al. (2024b) on LongBench. ThinK ($0.4$) indicates we prune the key cache channels with a pruning ratio of $\lambda=0.4$.
Method   Bit Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
KIVI 2/2 19.47 18.62 30.28 29.42 25.00 10.30 21.34 20.51 25.10 63.00 85.04 40.16 4.00 8.00 58.04 52.48 31.92
+ThinK (0.4) 2/2 19.46 19.01 30.52 28.79 25.78 9.53 22.11 20.66 25.73 63.00 84.62 41.54 3.50 7.00 56.51 48.92 31.77